OpenStudy (anonymous):

Use the Formula R = 6e12.77x, where x is the blood alcohol concentration and R, given in percent, is the risk of having a car accident. Solve: What is blood alcohol concentration corresponds to a 25% risk of an accident?

OpenStudy (jackellyn):

25 = 6e12.77x 25/6 = e12.77x ln(25/6) = ln(e12.77x) ln(25/6) = 12.77 x ln (25/6)/ 12.77 = x